# MotifDisco: Motif Causal Discovery For Time Series Motifs

### v1

@@ 7       $loss=-\sum_{i=1}^{b}\left(y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right)$
5 Uses of Learned Motif Causality
To illustrate the use of Motif Causality as a building block for other downstream tasks, in this section we integrate Motif Causality with three model use cases of Forecasting, Anomaly Detection and Clustering. We evaluate the performance of these use cases later in section 6.
@@ 5 Uses of Learned Motif Causality
Clustering. For clustering, there is typically a method to group clusters based on minimizing distances between each data point and the cluster centroid. These distances are computed using various distance measures such as Discrete Time Warping (DTW) (Sakoe & Chiba, 1978) for traces. To integrate MC with a basic clustering algorithm, we use the motif causality values as an additional distance metric. The intuition here is to add an element of causality to the clustering, such that as the algorithm learns, MC values within each cluster will be minimized and similar data points within the cluster should be motif causal of each other. An example is shown in Figure 11 in A.1: the MC value between the blue centroid and the blue data point to the right is high with 0.9, whereas the MC between the blue centroid and the green data point belonging to a different cluster is low at 0.11.
@@ 6 Evaluation
In this section we evaluate our causal discovery framework in terms of scalability and performance for three downstream task use cases: forecasting, anomaly detection and clustering.
Experimental Details. For all experiments, we use sets of single-day glucose traces randomly sampled across each month from January to December 2022, collected from Dexcom’s G6 Continuous Glucose Monitors (CGMs)  Akturk et al. (2021). Data was recorded every 5 minutes (total of 288 timepoints per trace) and each trace was aligned temporally from 00:00 to 23:59. All experiments were run 5 times with an 80/20% train/test split on an Intel Sky Lake 48 CPU VM with 192GB of RAM. Motifs were pulled from traces using the chopping method and we set $\gamma=0.7$ and $\lambda=0.5$.
Motif Causal Graphs. An example motif causal graph from the MotifDisco framework is shown in Figure 8. The edge color indicates the strength of the MC relationship; darker is stronger (closer to 1) while lighter is less strong (closer to 0). A.1 shows example MC between motifs of different $\tau$.
@@
Figure 9: Scalability varying 9(a) # of traces ($n$), 9(b) motif lengths ($\tau$) and 9(c) motif set sizes ($|\mathcal{M}|$).
Scalability. The scalability evaluation is shown in Figure 9, computing total time to train the MotifDisco causal discovery framework in seconds for 9(a) different numbers of traces ($n$), 9(b) motif lengths ($\tau$) and 9(c) motif set sizes ($|\mathcal{M}|$). In all experiments training was for 10 epochs. As to be expected, training time increases for larger $n$ and $|\mathcal{M}|$. Interestingly, time reduces as the $\tau$ increases, which may in part be because larger $\tau$s mean there are less total motifs (i.e., $|\mathcal{M}|$ is smaller).
Model Forecasting Anomaly Detection Clustering
Base $0.22\pm 0.05$ $0.83\pm 0.02$ $0.80\pm 0.03$ $0.67\pm 0.02$ $\mathbf{1.0\pm 0.0}$ $0.086\pm 1e{-6}$ $477.0\pm 1e{-3}$ $0.318\pm 1e{-2}$ $490.95\pm 6e{-2}$
Forecasting. For the forecasting task, we use a simple bidirectional LSTM as our base model. We set the sliding window size, $\tau$ and forecasting window to 6 (corresponding to 30 minutes of time). We train the causal discovery and forecasting models for 20 and 2000 epochs, respectively. The Root Mean Square Error (RMSE) is reported in Table 1; the MC model outperforms the base one.
Anomaly Detection. For this task, we build a simple autoencoder consisting of stacks of sequential dense layers. In the base model, an anomaly is detected if the Mean Absolute Error (MAE) of the reconstructed data (i.e., the trace fed through the encoder and then returned via the decoder) is less than a reconstruction threshold which we set as the standard deviation of the mean of the normal (non-anomalous) training data. In the MC-integrated version, we detect an anomaly if the predicted MC value between the previous time chunk (motif) and the current one is less than the edge prediction threshold $\theta$. We set the sliding window size and $\tau$ to 48, $\theta$ to 0.1 and train the causal discovery and anomaly models for 10 and 50 epochs, respectively. We compare the models by computing a set of classification metrics including accuracy, F1, Sensitivity and Specificity using the ground truth labeled anomalies. Results are reported in Table 1. For all metrics except Specificity, the MC-integrated model does better. Interestingly, the base model has better Specificity while the MC model has better Sensitivity - this indicates the MC model is better at identifying the anomalies (the true negatives) and the base is better at identifying the normal traces (the true positives). Example anomaly predictions are shown in Figure 10. Each graph plots the original input trace (in blue) and the reconstructed trace (in red, trace fed through the encoder + decoder) with the error between the two shaded in red. The window segments (dashed black lines) correspond to the motif and prediction window size and the model predictions are annotated in each window, colored by the correctness of the prediction. Green indicates a correct prediction, red indicates an incorrect one.
@@
Figure 10: Anomaly Detection for a trace with an obvious anomaly using the 10(a) base model and the model integrated with MC 10(b). False means an anomaly was predicted and the color corresponds to correctness of the prediction: green is a correct prediction, red is an incorrect one.
Clustering. We implement a simple K-Means clustering algorithm. In the base model we compute distances between the cluster centroids and other trace data points using Discrete Time Warping (DTW) (Sakoe & Chiba, 1978). In the MC-integrated version we compute the distances as DTW + MC. We set the number of clusters $k=3$, and $\tau=48$. Causal discovery and clustering models were trained for 20 epochs. We compute a set of clustering evaluation metrics including C-Index (Hubert & Levin, 1976), Sum of Squared Error (SSE) (Macqueen, 1967), Silhouette score (Rousseeuw, 1987), and Caliński Harabasz score (Caliński & Harabasz, 1974), reported in Table 1. Across all metrics the MC version does better, indicating adding an element of causality may help clustering.
7 Conclusion & Limitations
Appendix A Appendix
A.1 Additional Figures
Figure 11: Depiction of Clustering Model integrated with MC.
@@ A.1 Additional Figures
Figure 12: Example High Motif Causality values between different motif sizes.
@@
Figure 13: Example Low Motif Causality values between different motif sizes.

### v2

@@ 5 Uses of Learned Motif Causality
Clustering. For clustering, there is typically a method to group clusters based on minimizing distances between each data point and the cluster centroid. These distances are computed using various distance measures such as Discrete Time Warping (DTW) [39] for traces. To integrate MC with a basic clustering algorithm, we use the motif causality values as an additional distance metric. The intuition here is to add an element of causality to the clustering, such that as the algorithm learns, MC values within each cluster will be minimized and similar data points within the cluster should be motif causal of each other. An example is shown in Figure 12 in Appx. A.2.
@@ 6 Evaluation
As mentioned in section 2, existing Granger causal techniques are not directly comparable with MotifDisco due to their framework set-up or model assumptions; i.e., they formulate causality using multivariate statistical properties or temporal statistical measures using time lags which do not hold for motifs, or require labels or a known underlying causal structure which are not available for our health data streams (see Appx. A.3.) Therefore, in this section we evaluate our causal discovery framework in terms of scalability and performance for three downstream use cases: forecasting, anomaly detection and clustering.
Experimental Details.
We use three health stream datasets with different lengths and sampling frequencies: glucose traces from continuous glucose monitors [40], ECG traces and respiration (Resp) traces from the MIMIC-BP dataset [41]. All experiments were run 5 times with an 80/20% train/test split on an Intel Sky Lake 48 CPU VM with 192GB of RAM. Motifs were pulled from traces using the chopping method and we set $\gamma=0.7$ and $\lambda=0.5$. Additional details are available in Appx. A.3.
Motif Causal Graphs. An example motif causal graph from the MotifDisco framework is shown in Figure 7. The edge color indicates the strength of the MC relationship; darker is stronger (closer to 1) while lighter is less strong (closer to 0). A.3.4 shows example MC between motifs of different $\tau$.
Scalability. Scalability evaluation is shown in Figure 8 and Appx. A.3.2, computing total time to train the MotifDisco causal discovery framework in seconds for (a) different numbers of traces $n$, (b) motif lengths $\tau$ and (c) motif set sizes $|\mathcal{M}|$. Training was for 10 epochs in all experiments. As to be expected, training time increases for larger $n$ and $|\mathcal{M}|$. Interestingly, time reduces as the $\tau$ increases, which may in part be because larger $\tau$s mean there are less total motifs (i.e., $|\mathcal{M}|$ is smaller).
Forecasting. For the forecasting task, we use a simple bidirectional LSTM as our base model. We set the sliding window size, $\tau$ and forecasting window to 6 timesteps. We train the causal discovery and forecasting models for 20 and 2000 epochs, respectively. The Root Mean Square Error (RMSE) is reported in Table 1; the MC model outperforms the base one for all datasets.
Anomaly Detection. For this task, we build a simple autoencoder consisting of stacks of sequential dense layers. In the base model, an anomaly is detected if the Mean Absolute Error (MAE) of the reconstructed data (i.e., the trace fed through the encoder and then returned via the decoder) is less than a reconstruction threshold which we set as the standard deviation of the mean of the normal (non-anomalous) training data. In the MC-integrated version, we detect an anomaly if the predicted MC value between the previous time chunk (motif) and the current one is less than the edge prediction threshold $\theta$. We set the sliding window size and $\tau$ to 48, $\theta$ to 0.1 and train the causal discovery and anomaly models for 10 and 50 epochs, respectively. We compare the models by computing a set of classification metrics including accuracy, F1, Sensitivity and Specificity using the ground truth labeled anomalies. Results are reported in Table 1. For all metrics except Specificity, the MC-integrated model does better. Interestingly, the base model has better Specificity while the MC model has better Sensitivity - this indicates the MC model is better at identifying the anomalies (the true negatives) and the base is better at identifying the normal traces (the true positives). Example anomaly predictions are shown in Figure 15 in A.3.3.
Dataset Model Forecasting Anomaly Detection Clustering
Glucose Base $0.22\pm 0.05$ $0.83\pm 0.02$ $0.80\pm 0.03$ $0.67\pm 0.02$ $\mathbf{1.0\pm 0.0}$ $0.086\pm 1e{-6}$ $477.0\pm 1e{-3}$ $0.318\pm 1e{-2}$ $490.95\pm 6e{-2}$
ECG Base $0.30\pm 0.14$ $0.72\pm 0.10$ $0.74\pm 0.11$ $0.59\pm 0.13$ $0.70\pm 0.09$ $0.041\pm 6e{-2}$ $821.10\pm 2e{-3}$ $0.426\pm 3e{-3}$ $970.82\pm 4e{-3}$
MC $\mathbf{0.17\pm 0.12}$ $\mathbf{0.80\pm 0.11}$ $\mathbf{0.75\pm 0.09}$ $\mathbf{0.83\pm 0.09}$ $0.70\pm 0.12$ $\mathbf{0.040\pm 3e{-2}}$ $\mathbf{820.46\pm 1e{-2}}$ $\mathbf{0.428\pm 4e{-3}}$ $\mathbf{972.27\pm 2e{-4}}$
Resp Base $0.32\pm 0.11$ $0.65\pm 0.12$ $0.73\pm 0.13$ $0.31\pm 0.14$ $\mathbf{0.91\pm 0.15}$ $0.075\pm 2e{-3}$ $476.6\pm 3e{-4}$ $0.430\pm 3e{-2}$ $1006.12\pm 5e{-2}$
MC $\mathbf{0.26\pm 0.14}$ $\mathbf{0.71\pm 0.12}$ $\mathbf{0.74\pm 0.11}$ $\mathbf{0.65\pm 0.12}$ $0.86\pm 0.14$ $\mathbf{0.073\pm 5e{-3}}$ $\mathbf{476.5\pm 2e{-3}}$ $\mathbf{0.432\pm 4e{-2}}$ $\mathbf{1007.65\pm 3e{-4}}$
@@
Figure 8: Scalability for glucose traces varying 8(a) # of traces ($n$), 8(b) motif lengths ($\tau$) and 8(c) motif set sizes ($|\mathcal{M}|$). Results for the other data streams are available in Appx. A.3.2.
Clustering. We implement a simple K-Means clustering algorithm. In the base model we compute distances between the cluster centroids and other trace data points using Discrete Time Warping (DTW) [39]. In the MC-integrated version we compute the distances as DTW + MC. We set the number of clusters $k=3$, and $\tau=48$. Causal discovery and clustering models were trained for 20 epochs. We compute a set of clustering evaluation metrics including C-Index [42], Sum of Squared Error (SSE) [43], Silhouette score [44], and Caliński Harabasz score [45], reported in Table 1. Across all metrics the MC version does better, indicating adding an element of causality may help clustering.
7 Conclusion & Limitations
Appendix A Technical Appendices and Supplementary Material
A.1 Additional Introduction
Sample Traces. Sample real traces and motifs from glucose are available in Figure 1. Sample traces and motifs for ECG are shown in Figure 9 and for respiration (resp) are shown in Figure 10.
Figure 9: Real ECG Traces and Sample Motifs for $\tau=48$.
Figure 10: Real Respiration (Resp) Traces and Sample Motifs for $\tau=48$.
A.2 Additional Methodology
Extracting Motifs. We assume there is a consistent, conclusive way to pull motifs from the traces. Motif discovery is outside the scope of this work; discussion about motif extraction methods and artifacts is outside the scope of the paper since the problem of how to extract motifs has been widely solved (e.g., there are many nice methods to pull out motifs in an intelligent manner, given some criteria, and specific to different disciplines, such as  [11, 25, 26]). Our framework is amenable to any motif extraction method, and, as evidenced by our results, works even when using simple or heuristic motif extraction methods, such as chopping.
(a) Large Uncertainty Reduction
(b) Small Uncertainty Reduction
Figure 11: Depiction of Conditional Entropy $H(\mu)$ computation using the histogram method: when computing MC for $\mu^{i}$, 11(a) $\mu^{j}$ results in a large uncertainty reduction; 11(b) $\mu^{z}$ has a small reduction.
MC Computation in Link Predictor. To implement the MC computation, we adapt existing Transfer Entropy libraries [38] and compute the conditional entropy function (i.e., $H(\mu)$, Shannon’s or Rényi entropy) using the histogram method. A simplified depiction is shown in Figure 11. Essentially, motif time series values are binned into a histogram. Distributions between motifs can be compared to determine how much uncertainty about future predictions of the motif is reduced in the distribution. For example, in 11(a) when computing MC for $\mu^{i}$, $\mu^{j}$ covers a larger distribution, resulting in a large reduction in uncertainty and higher motif causality, whereas 11(b) $\mu^{z}$ covers hardly any new distribution compared to the status quo ($H(\mu^{i}|\mathcal{K})$), resulting in a small reduction in uncertainty and low causality for $\mu^{z}$.
Clustering Use Case. To integrate Motif Causality with a basic clustering algorithm, we use the motif causality values as an additional distance metric. The intuition is to add an element of causality to the clustering, such that as the algorithm learns, MC values within each cluster will be minimized and similar data points within the cluster should be motif causal of each other. An example is shown in Figure 12: the MC value between the blue centroid and the blue data point to the right is high with 0.9, whereas the MC between the blue centroid and the green data point belonging to a different cluster is low at 0.11.
Figure 12: Depiction of Clustering Model integrated with MC.
A.3 Additional Evaluation
As mentioned in section 2, there is no prior work on quantifying causality amongst motifs, and we cannot compare with existing Granger causal techniques because their framework set-up or model assumptions do not hold. Specifically, they have one or more of the following issues:
  • Compose relations between only two time series, and cannot find relationships amongst a set of motifs or traces [4, 28].
  • Formulate causality based on various statistical properties amongst variables, such as variable-based correlation and density-, classification- and prediction-based error measures computed between multivariate features [18, 28, 29]. These assumptions do not hold for univariate time series motifs.
  • Compose causal relationships over time series using repeated statistical measures or temporal dynamics, e.g., things like learning implicit repeated patterns, or computing sufficiency or faithfulness hypotheses over lags across the full time series [12, 15, 16, 17, 19, 30, 31]. These methods do not work for motifs because they are short time series (so there would be no repeated patterns, nor will statistical hypotheses like sufficiency or faithfulness hold since they are not evaluating multiple time lags repeated over a trace).
  • Require labels or a known underlying causal structure to guide model training [4, 18, 19], which is not available for our data and many similar event-based data streams.
Therefore, for our evaluation in section 6, we evaluate our causal discovery framework in terms of scalability and performance for 3 downstream use cases: forecasting, anomaly detection and clustering.
A.3.1 Dataset Details
An overview of characteristics for each dataset is shown in Table 2. The glucose traces are sets of single-day glucose traces randomly sampled across each month from January to December 2022, collected from Dexcom’s G6 Continuous Glucose Monitors (CGMs) [40]. Data was recorded every 5 minutes, resulting in a total of 288 timepoints per trace. Each trace was aligned temporally from 00:00 to 23:59. The ECG and Respiration (resp) traces come from MIMIC-BP [41], a curated dataset collecting common biomedical signals. We use data from a single segment. There are a total of 1523 patients, with data collected for a total of 3750 timepoints, sampled at frequencies of 1 millisecond for ECG and 1 second for Resp.
Table 2: Dataset Characteristics.
Dataset # Patients Total # Timepoints Sampling Frequency
Glucose [40] 300000 288 5 minutes
ECG [41] 1523 3750 1 millisecond
Resp [41] 1523 3750 1 second
A.3.2 Additional Scalability Results
Due to space constraints, we report the results for the MIMIC-BP dataset streams of ECG and Respiration (Resp) in Figure 13 and Figure 14 respectively here. As found with the glucose traces, training time increases for larger $n$ and $|\mathcal{M}|$. Time also reduces as the $\tau$ increases, which may in part be because larger $\tau$s mean there are less total motifs (i.e., $|\mathcal{M}|$ is smaller). Overall, training time for the Resp dataset takes the longest, perhaps because the Resp traces are the more variable, resulting in more variable (i.e., more noisy) motifs.
(a) $n$ vs. Training Time
(b) $\tau$ vs. Training Time
(c) $|\mathcal{M}|$ vs. Training Time
Figure 13: Scalability for ECG traces varying 13(a) # of traces ($n$), 13(b) motif lengths ($\tau$) and 13(c) motif set sizes ($|\mathcal{M}|$).
(a) $n$ vs. Training Time
(b) $\tau$ vs. Training Time
(c) $|\mathcal{M}|$ vs. Training Time
Figure 14: Scalability for respiration traces varying 14(a) # of traces ($n$), 14(b) motif lengths ($\tau$) and 14(c) motif set sizes ($|\mathcal{M}|$).
A.3.3 Additional Use Case Results
Anomaly Detection. Example anomaly predictions are shown in Figure 15. Each graph plots the original input trace (in blue) and the reconstructed trace (in red, trace fed through the encoder + decoder) with the error between the two shaded in red. The window segments (dashed black lines) correspond to the motif and prediction window size and the model predictions are annotated in each window, colored by the correctness of the prediction. Green indicates a correct prediction, red indicates an incorrect one.
@@ A.3.3 Additional Use Case Results
Figure 15: Anomaly Detection for a trace with an obvious anomaly using the 15(a) base model and the model integrated with MC 15(b). False means an anomaly was predicted and the color corresponds to correctness of the prediction: green is a correct prediction, red is an incorrect one.
A.3.4 Example Motif Causality Between Motifs
Learned motif causality values between sample motifs for glucose are illustrated in Figure 16 and Figure 17.
A.4 Additional Conclusion
Broader Impacts & Safeguards. The core contribution of this project is the development of MotifDisco, the first causal discovery framework to infer Motif Causality amongst time series motifs. In terms of broader impacts, MotifDisco may facilitate the development of advanced, high performing technologies for event-based time series by providing a new method to learn and quantify relationships amongst motifs. Since this project presents a foundational research framework that can be applied in many diverse settings and scenarios, discussion of deployments is outside the scope of this work. However, the authors note that any deployment using this framework may need to integrate the necessary measures and safeguards to reduce the risk of possible societal harms (particularly for applications in medical settings), such as data release and consent frameworks, and strict privacy and security measures.
@@ A.4 Additional Conclusion
Figure 16: Example High Motif Causality values between different motif sizes for Glucose.
@@
Figure 17: Example Low Motif Causality values between different motif sizes for Glucose.